When considering the set of all the natural numbers (ℕ), show whether the mathematical operations of addition, subtraction, multiplication and division are: (a) Associative (b) Commutative.

1
Expert's answer
2021-11-26T11:49:20-0500

(a) Since "a+(b+c)=a+b+c=(a+b)+c" for any "a,b,c\in\N," we conclude that the operation of addition is associative on the set "\N" of natural numbers.

Since "a\cdot(b\cdot c)=a\cdot b\cdot c=(a\cdot b)\cdot c" for any "a,b,c\in\N," we conclude that the operation of multiplication is associative on the set "\N" of natural numbers.

Since "4-(2-1)=3\ne 1=(4-2)-1," we conclude that the operation of subtraction is not associative on the set "\N" of natural numbers.

Since "8:(4:2)=4\ne 1=(8:4):2," we conclude that the operation of division is not associative on the set "\N" of natural numbers.


(b) Since "a+b=b+a" for any "a,b\in\N," we conclude that the operation of addition is commutative on the set "\N" of natural numbers.

Since "a\cdot b=b\cdot a" for any "a,b\in\N," we conclude that the operation of multiplication is commutative on the set "\N" of natural numbers.

Since "2-1=1\ne -1=1-2," we conclude that the operation of subtraction is not commutative on the set "\N" of natural numbers.

Since "8:4=2\ne 0.5=4:8," we conclude that the operation of division is not commutative on the set "\N" of natural numbers.